Two years ago a man was six times as old as his daughter. In 18 years he will be twice as old as his daughter. Determine their present ages.
Answer by ankor@dixie-net.com(22740)

let m = man's present age
let d - daughter's
:
Write an equation for each statement, simplify as much as possible
:
Two years ago a man was six times as old as his daughter.
m- 2 = 6(d-2)
m - 2 = 6d - 12
m = 6d - 12 + 2
m = 6d - 10
:
In 18 years he will be twice as old as his daughter.
m + 18 = 2(d+18)
m + 18 = 2d + 36
m = 2d + 36 - 18
m = d + 18
From the 1st equation, replace m with (6d-10)
6d - 10 = 2d + 18
6d - 2d = 18 + 10
4d = 28
d = 28/4
d = 7 yrs is the daughter's age
Find Dad's age
m = 6(7) - 10
m = 42 - 10
m = 32 yrs is Dad's age
:
:
Check solutions in the statement:
"In 18 years he will be twice as old as his daughter."
32 + 18 = 2(7 + 18)
50 = 2(25)